OpenStudy (anonymous):

$15,070 with a 7% sales tax?

sam (.sam.):

7/100*15070=1054.9 Extra to pay she has 1500, 1500-1054.9=445.1 extra

Directrix (directrix):

15 070 * 1.07 = 16 124.90

OpenStudy (kymber):

First you can convert 7% to a decimal, by moving the decimal point, it's .07 and then you can multiply .07 by $15,070 which is $1,054.9 Then adding them together, which is $16,124.9 Don't forget the dollar sign! Hope this helps :D